Food is sufficient for 100men for 60days. For how many days the food is sufficient for 500men?

Grade:

2 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

Required no of days=300 days...

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi

Sudheesh Singanamalla
13 years ago

100 / 60 = 500 / X

1/6 = 5/X

X = 30 days !

Food is sufficient for 500 men for 30 days